OpenStudy (anonymous):

Find the equation of the line through the given points. (3, 4) and (1, -2)

OpenStudy (anonymous):

y=3x-5

OpenStudy (ybarrap):

slope = (-2 - 4) / (1 - 3) = -6/-2 = 3 y = 3x + b using (3,4) 4 = 3(3) + b b = 4 - 3*3 = 4 - 9 = -5 Ans. y = 3x - 5